(1 point) a. Find the most general real-valued solution to the linear system of differential equations = 1 x1 (1) = C1 + C2I AM REALLY STRUGGLING ON THIS PROBLEM PLEASE HELP ME CORRECT AND NEAT WORK IS MUCH APPRECIATED THANKS
Answer #1

solution: ca) *- []* o per me (* *]p] btw 21 (3) -6 first we find eigen value of - let a be an eigen value of tour then 1 IA-=) HF-22 de 24=di ui] xit Cgetzt, a) - diru -482=0 and 24 - 21x2=0. -) aixo - (911²x220 =) gix + 4x2 = 0 =) dixi=-482 =) ix =24 (1) (top] сі е total ai] + ca 2-ai X2 (4) fai & staisty +c2 to-aut. -di 66-21) fot&ilt 210) = dig66+ (6+2) (-6-ai)t + (-ai(6) Result: Consider the linear system da at ax+by 3-c را) dy catdy dt of the form dizide and if both chasectecistic sols ofPlease give a rating if it was helpful.. Thank you